Let X be a continuous random variable distributed by N(5,0.2) find,

A) P(5.9 < X < 5.9)

B) P( 4.6 =< X =< 5.3)

C) P(5.6=< X < 5.7)

D) P( 5.1< X =< 5.7)
Answer #1

Solution :

Given that ,

mean = = 5

standard deviation = = 0.2

A) P( 5.9 < x < 5.9) = P[( 5.9 - 5)/ 0.2) < (x - ) /  < (5.9 - 5) / 0.2) ]

= P( 4.5 < z < 4.5)

= P(z < 4.5) - P(z < 4.5)

Using z table

= 1 - 1

= 0

B) P( 4.6 x 5.3 )

= P[(4.6 - 5 / 0.2) (x - ) / ( 5.3 - 5 / 0.2) ]

= P(-2.00   z 1.50 )

= P(z 1.50 ) - P(z -2.00)

Using z table

= 0.9332 - 0.0228

= 0.9104

C) P( 5.6 x < 5.7 )

= P[(5.6 - 5 )/0.2 ) (x - ) /  < ( 5.7 - 5) / 0.2) ]

= P(3.00 < z < 3.50)

= P(z < 3.50 ) - P(z < 3.00 )

Using z table

= 0.9999 - 0.9987

= 0.0012

D) P(5.1 < x 5.7 )

= P[(5.1 - 5 )/0.2 ) (x - ) /  < ( 5.7 - 5) / 0.2) ]

= P(0.50 < z < 3.50)

= P(z < 3.50 ) - P(z < 0.50 )

Using z table

= 0.9999 - 0.6915

= 0.3084